OpenStudy (fools101):

~Please help!~ Solve the following equation for x. 4x+7+2x=3x+x-13 A) x=10 B) x=2 C) x=-10 D) x=-3

OpenStudy (lawliet):

answer is c)x=-10

OpenStudy (anonymous):

use the distributive property for example 4x+2x=(4+2)x=6x just continue simplifying the problem. move all the variables to one side and divide both sides by the coefficient in order to get the answer

OpenStudy (fools101):

AWWW thank u Anhhuyalex and Lawliet ^ ^

OpenStudy (anonymous):

@anhhuyalex Do people use the distributive property to add terms? Because that is kind of pointless.

OpenStudy (fools101):

~one more help please!~ Which angle is an obtuse angle? A) 67° B) 203° C) 154° D) 90°

OpenStudy (anonymous):

no, we just intuitively apply the distribute property. 4x+2x=6x. You just know that because it is so obvious

OpenStudy (anonymous):

oh, and @fools101, an obtuse angle is an angle whose measure is between 90 and 180

OpenStudy (lawliet):

A)Acute B)Reflex C)Obtuse D)Right

OpenStudy (fools101):

C?

OpenStudy (fools101):

AWWW kk it would be C) 154° ^ ^... kk i got it thank u guys for the help...